OpenStudy (anonymous):

What is the slope of the line perpendicular to -2y=3x+5 ?

OpenStudy (anonymous):

slope of the given line is m=-3/2 slope of perpendicular is =-1/m=-1/(-3/2)=2/3

OpenStudy (klimenkov):

\[2/3.\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@klimenkov Is wrong btw....let this be a lesson that we should not take answers only without explanation

OpenStudy (anonymous):

-2y=3x+5 =y=-3/2x-5/3 the slope of this line equation is -3/2. since the product of slops of perpendicular line is =-1, we get =z(-3/2)=-1 =z=2/3 so the slope of the perpendicular line is 2/3. ..

OpenStudy (anonymous):

@Romero its ok i get it now

OpenStudy (klimenkov):

sorry, i'm just a novice

OpenStudy (anonymous):

@klimenkov dont be sorry! You helped me out